OpenStudy (erinweeks):

The following piecewise function gives the tax owed, T(x), by a single taxpayer on a taxable income of x dollars.

OpenStudy (erinweeks):

T(x) = (i) Determine whether T is continuous at 6061. (ii) Determine whether T is continuous at 32,473. (iii) If T had discontinuities, use one of these discontinuities to describe a situation where it might be advantageous to earn less money in taxable income.

OpenStudy (erinweeks):

T(x) =

OpenStudy (anonymous):

Oh, shhiii...take mushrooms lol. That is a huge function haha Lets see: \[T(x)=\left\{\eqalign{ &0.10x; &x\in(0,6061]\\ &606.10+0.18(x-6061); & x\in (6061,32473]\\ &5360.26+0.26(x-32473); &x\in (32473,72784] \\ &15841.12+0.29(x-72784); &x\in (72784, 149897] \\ &38203.89+0.32(x-149897); &x\in (149897,325127]\\ &94277.49+0.36(x-325127); &x\in(325127,\infty) \\ }\right.\]

OpenStudy (erinweeks):

wow im already confused. -_-

OpenStudy (erinweeks):

@ganeshie8 may you explain?

OpenStudy (anonymous):

That's the same function. lol

OpenStudy (erinweeks):

okay but what do i do from here

OpenStudy (anonymous):

Erm...is this a calculus question?

ganeshie8 (ganeshie8):

say, you're working and earning X amount per year

ganeshie8 (ganeshie8):

ur salary is X dollars

OpenStudy (erinweeks):

yea a pre calc question ... & okay

ganeshie8 (ganeshie8):

you will have to pay tax on ur salary X, here, how much % u pay as tax depends on how much u earning

OpenStudy (erinweeks):

okay!

ganeshie8 (ganeshie8):

look at first line, if u earn \(\le 6061\), then u will have to pay a tax of 10%

OpenStudy (anonymous):

I'm leaving this to @ganeshie8 haha

ganeshie8 (ganeshie8):

okie.. ive time :) @KeithAfasCalcLover

OpenStudy (erinweeks):

okay

ganeshie8 (ganeshie8):

first understand all the lines, it simply tells us, how much u have to pay as tax. next, we can work the discontinuity part

OpenStudy (erinweeks):

hold on how did you get 10%

ganeshie8 (ganeshie8):

0.10x is same as 10% of x

OpenStudy (erinweeks):

okay so basically all of them are 10 %?

ganeshie8 (ganeshie8):

nope, second line is 18% third line is 26% fourth line is 29% ..... last line is 36%

ganeshie8 (ganeshie8):

more u earn, more the government sucks tax money from u

OpenStudy (erinweeks):

okay and five line is 32 ! gotcha !!

ganeshie8 (ganeshie8):

Yup ! lets plot this piecewise function and see, where its continuous and discontinuous

ganeshie8 (ganeshie8):

do u have geogebra ?

OpenStudy (erinweeks):

i used to but i started messing up and i cant re download it idk why?

ganeshie8 (ganeshie8):

okie let me plot it and attach a screenshot

OpenStudy (erinweeks):

thank you

ganeshie8 (ganeshie8):

okay, graphing is bit tedious lets do some algebra quick

ganeshie8 (ganeshie8):

(i) Determine whether T is continuous at 6061.

OpenStudy (erinweeks):

okay lol

ganeshie8 (ganeshie8):

\( T(x)=\left\{\eqalign{ &0.10x; &x\in(0,6061]\\ &606.10+0.18(x-6061); & x\in (6061,32473]\\ &5360.26+0.26(x-32473); &x\in (32473,72784] \\ &15841.12+0.29(x-72784); &x\in (72784, 149897] \\ &38203.89+0.32(x-149897); &x\in (149897,325127]\\ &94277.49+0.36(x-325127); &x\in(325127,\infty) \\ }\right. \)

ganeshie8 (ganeshie8):

when x = 6061, when u go from left :- T(x) = .10(6061) = 606.10 when u go from right :- T(x) = 606.10 + .18(6061-6061) = 606.10+ 0 = 606.10

ganeshie8 (ganeshie8):

left side value = right side value so function is continuous at x = 6061

ganeshie8 (ganeshie8):

does that make some sense

OpenStudy (erinweeks):

made a little sense....

ganeshie8 (ganeshie8):

(ii) Determine whether T is continuous at 32,473.

ganeshie8 (ganeshie8):

do this same way, calculate left side & right side values at x = 32473

OpenStudy (erinweeks):

holdd onn so for (i) i dont get what im suppose to put???

ganeshie8 (ganeshie8):

put below :- "T(x) is continuous at x=6061 because both left side and right side values are same : 606.10 "

ganeshie8 (ganeshie8):

??

OpenStudy (erinweeks):

sorrry okay gotcha ... was confused...

ganeshie8 (ganeshie8):

its okay, once u do part ii, u wil get some confidence im sure

ganeshie8 (ganeshie8):

See the tax table once

ganeshie8 (ganeshie8):

\( T(x)=\left\{\eqalign{ &0.10x; &x\in(0,6061]\\ &606.10+0.18(x-6061); & x\in (6061,32473]\\ &5360.26+0.26(x-32473); &x\in (32473,72784] \\ &15841.12+0.29(x-72784); &x\in (72784, 149897] \\ &38203.89+0.32(x-149897); &x\in (149897,325127]\\ &94277.49+0.36(x-325127); &x\in(325127,\infty) \\ }\right. \)

ganeshie8 (ganeshie8):

for x = 32473, left side function is SECOND LINE right side function is THIRD LINE

OpenStudy (erinweeks):

so hold on i did .24(32473) = 8442.98 i got ?

ganeshie8 (ganeshie8):

evaluate both left side & right side functions for x = 32473, and see if u get the same value

ganeshie8 (ganeshie8):

you need to use SECOND LINE for left side function

OpenStudy (erinweeks):

okay did i do it right for the left side??

ganeshie8 (ganeshie8):

no, from where u got .24 ?

ganeshie8 (ganeshie8):

left side function :- 606.10+0.18(x−6061)

ganeshie8 (ganeshie8):

right side function :- 5360.26+0.26(x−32473)

ganeshie8 (ganeshie8):

u just need to put x = 32473 above and see if u get same value

OpenStudy (erinweeks):

first one i got 5360.26 second one i got 5360.26 so yea its equal..

ganeshie8 (ganeshie8):

perfect ! since left side value = right side value, at x = 32473, T is continuous at x = 32473

ganeshie8 (ganeshie8):

getting ?

OpenStudy (erinweeks):

yea im getting it a little more... but now how do we do (iii)

ganeshie8 (ganeshie8):

for iii) T is continuous at every point, it has no discontinuities. So earning less money for saving tax is not a good idea.

ganeshie8 (ganeshie8):

in other words, wid the given piecewise function, since it is continuous, it is never advantageous to earn less money in taxable income

OpenStudy (erinweeks):

okay thankk you soo muchhh!

ganeshie8 (ganeshie8):

np :) iii part is just there to distract you. It really doesnt help earning less money here. cuz this piecewise function is fully continuous in its range